OpenStudy (kva1992):

How would you solve 4cos^2x -4cosx=2. I tried the quadratic but dont think i did it properly. answers are 111.5 degrees and 248.5 degrees

OpenStudy (kva1992):

i first tried to factor a two after i moved the 2 over so i ended up with 2(2cos^2-2cosx-1)=0

OpenStudy (campbell_st):

this is an equation that can be solved as a quadratic divide every term by 2 and you get \[2\cos^2(x) -2\cos(x) -1 = 0\] now solve using the general quadratic formula a = 2, b = 2 and c = -1 hope it helps

OpenStudy (kva1992):

hmmm i think i did that but ill do it again

OpenStudy (kva1992):

i keep getting stuck at (cosx= 2 +or - 2 squareroot 3) / 4

OpenStudy (campbell_st):

well you get \[\frac{2 \pm \sqrt{4 - 4 \times 2 \times -1}}{2 \times 2} = \frac{2 \pm \sqrt{12}}{4}\] which simplifies to \[\cos(x) = \frac{1 \pm \sqrt{3}}{2}\]

OpenStudy (kva1992):

yup i got that

OpenStudy (kva1992):

how how do i find the degrees?

OpenStudy (kva1992):

inverse cos?

OpenStudy (campbell_st):

just go \[x = \cos^{-1}(\frac{1 + \sqrt{3}}{2})\] then the other one with a negative infront of the square root

OpenStudy (kva1992):

oh okie cool thanks

OpenStudy (kva1992):

really appreciate it so i was on the right track i just kept trying to find the + when that isnt possible because it exceeds the range.

OpenStudy (kva1992):

alright so my second question would be how did they find 248.5 degrees?

OpenStudy (campbell_st):

well that's correct... there is an error with the question since to solution is larger than 1, there is no angle... cos(x) is always between -1 and 1

OpenStudy (campbell_st):

yep... so its only the 2nd solution (1 - sqrt{3})/2 so the angle is in the 2nd and 3rd quadrant...

OpenStudy (kva1992):

but the answer sheet says its both 111.5 and 248.5

OpenStudy (kva1992):

and i have no idea how they figured out 248.5

OpenStudy (campbell_st):

thats correct... \[\frac{1 - \sqrt{3}}{2} = -0.366025\] so this is what I do... find the angle for the positive value \[x = \cos^{-1}(0.366025)\] so the angle is 68.63 degrees in the 2nd quadrant the angle is 180 - 68.63 in the 3rd quadrant the angle is 180 + 68.63 hope this helps

OpenStudy (kva1992):

oh yaaaa i completely forgot you arent supposed to find the negative of an inverse function duh

OpenStudy (campbell_st):

the cos ratio is negative in the 2nd and 3rd quadrants... 2nd is always 180 - angle 3rd is always 180 + angle

OpenStudy (kva1992):

so if i was trying to find the sin it would be in either q1 or q4 right?

OpenStudy (campbell_st):

well all ratios are positive in the 1st quadrant 2nd only sin is positive 3rd only tan is positive 4th only cos is positive and the angles are 2nd 180 - angle 3rd 180 + angle 4th 360 - angle hope it makes sense

OpenStudy (campbell_st):

well if sin was positive its 1st and 2nd...

OpenStudy (kva1992):

but doesnt the range influence how it works? like i remember with sine it goes from pi/2 to -pi/2

OpenStudy (kva1992):

or am i thinking of something completely different?

OpenStudy (ikram002p):

still stuck with this ?

OpenStudy (kva1992):

na figured the answer out just clearing up a question is all

OpenStudy (ikram002p):

^^

OpenStudy (campbell_st):

so if the range is from -pi/2 to pi/2 you are working in the 4th and 1st quadrant... sin is positive in the 1st and negative in the 4th

OpenStudy (campbell_st):

here is a very detailed cheat sheet on trig it may halp http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet_Reduced.pdf

OpenStudy (kva1992):

alright cool ill look into thanks really appreciate your time and help :)